OpenStudy (anonymous):

what is the answer to... 18+-17 ill give a medal

OpenStudy (anonymous):

1

OpenStudy (anonymous):

18+-17=1 18-17=1 you can move the add away

OpenStudy (anonymous):

1!

OpenStudy (anonymous):

just 1oh its 18+-17

OpenStudy (anonymous):

yep is that easy +- = - - - = + remember that

OpenStudy (anonymous):

i have 3 more questions can you help

OpenStudy (anonymous):

sure i guess

OpenStudy (anonymous):

ok well -12+4

OpenStudy (anonymous):

-12+4 = 12 - 4 and 12 - 4 = 8

OpenStudy (anonymous):

so 8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can switch the - and + For Exm: -1+2=? switch the -in 1 to + in 2 = +1-2 you can ignore the +1 is positive so you get 1-2= -1

OpenStudy (anonymous):

ok here we go -9+-2=

OpenStudy (anonymous):

is still the same if there is 3 marks is just like add or subtract -9+-2= 9+2= 11 but is - adding so is -11

OpenStudy (anonymous):

ok well thank you so much i hope you got the umm messange i sent you!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

so can you tutor me

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

just ask me if you need help

OpenStudy (anonymous):

ok well ill messange you when ever ur not busy that way i can now what day you can help me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@bRidingHoo isn't -12+4 = -8

OpenStudy (anonymous):

i didnt press the negative sign? ..

OpenStudy (anonymous):

i missed that oops

OpenStudy (anonymous):

cool, that happens to me to